A car welghing 9800 n is moving with a speed of 40 km/h.
on the application of the brakes it comes to rest after travel ling a distance of 50 metres. calculate the average retarding force.
(ans. 1234.57 ni
1
Expert's answer
2017-10-25T15:16:07-0400
The mass of the car:
m=W/g=9800/9.8=1000 kg.
The speed of car:
v=40 1000/3600=100/9 m/s
v^2-0^2=2ad
a=v^2/2d.
The average retarding force:
F=ma=m v^2/2d=1000 (100/9)^2/2(50) =1234.57 N.